# Re-Reading Improves Reasoning in Large Language Models

### v2

@@ Abstract
[…]
@@ 2.2 Re-Reading (Re2) Improves Reasoning
Re2 Prompting
Q: {Input Query} Read the question again: {Input Query} # Thought-eliciting prompt (e.g.,‘‘Let’s think step by step") #
@@ 3 Experiments
Table 1: Results on arithmetic reasoning benchmarks. $*{}^{*}$ denotes that Vanilla is even superior to CoT prompting.
@@
Table 2: Results on commonsense and symbolic reasoning benchmarks. $*{}^{*}$ denotes that Vanilla is even superior to CoT prompting.
@@ 3.3 Evaluation Results
When applied to ChatGPT, Re2 exhibits consistent improvement on most datasets, except for a slight drop in performance on a few datasets, e.g., AQUA and MultiArith, when using Vanilla+Re2. This exception could be due to ChatGPT’s exposure to these datasets with CoT outputs during instruction fine-tuning (IFT) Chen et al. (2023). On such datasets, ChatGPT with Vanilla still produces CoT-like output (see examples in Appendix E) and even outperforms ChatGPT with CoT (as indicated by the $*{}^{*}$ results in Tables 1 and 2). Chen et al. (2023) obtained similar experimental results and suggested that this occurs because ChatGPT may have been exposed to these task datasets containing CoT explanations without explicit prompting. Therefore, additional explicit instructions, like CoT or Re2, might disrupt this learned pattern in ChatGPT, possibly leading to decreased performance. Nonetheless, on some datasets like SVAMP, ASDIV, CSQA, and Date, Re2 still manages to improve the baseline Vanilla prompting. Moreover, in datasets where CoT prompting normally surpasses Vanilla prompting, such as GSM, StrategyQA, and Coin, Re2 significantly enhances Vanilla prompting ($\uparrow$ 4.63 on StrategyQA and $\uparrow$ 5.20 on the Coin dataset). Overall, our Re2 method still achieves improvements in 71% of the experiments on ChatGPT. More examples from the experiment results can be found in Appendix E.
@@ 3.4 Discussions
P0 Q: {question} #Answer format instruction# A: Let’s think step by step. 77.79 78.77
P1 Q: {question} Read the question again: {question} #Answer format instruction# A: Let’s think step by step. 79.45 80.59
P2 Q: {question} Q: {question} #Answer format instruction# A: Let’s think step by step. 78.09 79.38
P3 Q: {question} A: Let’s think step by step. Read the question again: {question} #Answer format instruction# A: Let’s think step by step. 79.08 80.36
P4 Q: {question} A: Let’s think step by step. Q: {question} #Answer format instruction# A: Let’s think step by step. 78.09 79.38
@@ Appendix E More Cases
Vanilla Q: {question} #Answer format instruction# A:
Vanilla+Re2 Q: {question} Read the question again: {question} #Answer format instruction# A:
CoT Q: {question} #Answer format instruction# A: Let’s think step by step.
CoT+Re2 Q: {question} Read the question again: {question} #Answer format instruction# A: Let’s think step by step.
PS Q: {question} #Answer format instruction# A: Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan, solve the problem step by step, and give the ultimate answer. Please explicitly generate the mentioned process: [Problem Understanding], [Plan], [Solving/Calculations], [Answer]. in your response.
PS+Re2 Q: {question} Read the question again: {question} #Answer format instruction# A: Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan, solve the problem step by step, and give the ultimate answer. Please explicitly generate the mentioned process: [Problem Understanding], [Plan], [Solving/Calculations], [Answer]. in your response.
@@
PAL #!/bin/python3 import math import numpy as np import statistics import sympy as sp ########### Task Instruction ############## # You will write python program to solve math problems. # You will only write code blocks. # Please generate your code block in ‘def solution()‘ function, and thus it can be executed by python interpreter. You don’t need to call ‘solution()‘ function because it will be called by the system. # The concrete format of ‘solution()‘ is as follows: # def solution(): # """<question>""" # <your code> # result = <your result> # return result ######################################### # Q: {question} # Your defined "solution()" function with comments here.
PAL+Re2 #!/bin/python3 import math import numpy as np import statistics import sympy as sp ########### Task Instruction ############## # You will write python program to solve math problems. # You will only write code blocks. # Please generate your code block in ‘def solution()‘ function, and thus it can be executed by python interpreter. You don’t need to call ‘solution()‘ function because it will be called by the system. # The concrete format of ‘solution()‘ is as follows: # def solution(): # """<question>""" # <your code> # result = <your result> # return result ######################################### # Q: {question} # Read the question again: {question} # Your defined "solution()" function with comments here.
@@
GSM Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
SVAMP Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
ASDIV Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
@@
MultiArith Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
SingleEQ Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
AddSub Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$, at the end of your response.
@@
StrategyQA Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
@@
Date Understanding Your answer should be a date, in the format of $𝑀𝑀𝐷𝐷𝑌𝑌𝑌𝑌$, e.g. 05/01/20220501202205 / 01 / 2022.
Coin Flip Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
@@
First, let’s calculate how much feed Wendi gives her chickens in the morning. She has 20 chickens, and she gives each chicken 3 cups of feed, so in total, she gives them 20 x 3 = 60 cups of feed in the morning.
Next, let’s calculate how much feed Wendi gives her chickens in the afternoon. Again, she has 20 chickens, and she gives each chicken 3 cups of feed, so in total, she gives them 20 x 3 = 60 cups of feed in the afternoon.
To find out how much feed she needs to give her chickens in the final meal of the day, we need to subtract the total feed given in the morning and afternoon from the total feed needed for the day. The total feed needed for the day for 20 chickens is 20 x 3 x 3 = 180 cups of feed.
So, the feed needed in the final meal of the day is 180 - 60 - 60 = 60 cups of feed.
Therefore, Wendi needs to give her chickens 60 cups of feed in the final meal of the day.
Answer: 606060
@@
First, we need to determine the total amount of feed Wendi gives her chickens in the morning and afternoon. In the morning, she gives 15 cups of feed, and in the afternoon, she gives another 25 cups. So, the total amount of feed given in the morning and afternoon is 15 + 25 = 40 cups.
Next, we need to determine the total amount of feed needed for Wendi’s flock of 20 chickens. Since each chicken requires 3 cups of feed, we multiply 3 by 20 to get 60 cups.
Finally, we subtract the total amount of feed given in the morning and afternoon from the total amount needed. 60 - 40 = 20
Therefore, Wendi needs to give her chickens 20 cups of feed in the final meal of the day.
Answer: 202020
@@
First, let’s calculate the cost of the movies Colby has already seen. He has seen 5 movies on a Friday or Saturday, which cost \$10 each, so that is a total of 5 * \$10 = \$50. He has also seen 8 movies on other days, which cost \$7 each, so that is a total of 8 * \$7 = \$56. Therefore, the total cost of the movies he has already seen is \$50 + \$56 = \$106.
Next, let’s calculate the cost of the popcorn and candy he wants to buy on the last day of the month. The popcorn costs \$8 and the candy costs \$2, so the total cost of the snacks is \$8 + \$2 = \$10.
Now, let’s subtract the cost of the movies Colby has already seen and the cost of the snacks from the \$150 his parents gave him. \$150 - \$106 - \$10 = \$34.
Since each movie on a Friday or Saturday costs \$10 and Colby has \$34 left, he can see a maximum of \$34 / \$10 = 3 movies on a Friday or Saturday.
In conclusion, Colby can see a total of 3 + 5 = 8 movies in the month.
888
@@
First, let’s calculate the total amount Colby has spent so far:
For the 5 movies on Fridays and Saturdays, he spent 5 x \$10 = \$50. For the 8 movies on other days, he spent 8 x \$7 = \$56. For the 2 tubs of popcorn, he spent 2 x \$8 = \$16. For the 4 boxes of candy, he spent 4 x \$2 = \$8.
Adding all these expenses together, Colby has spent \$50 + \$56 + \$16 + \$8 = \$130.
Now, let’s calculate how much money he has left from the \$150 his parents gave him:
Colby started with \$150 - \$130 = \$20.
Since it’s the last day of the month and it’s a Friday, he wants to buy a popcorn (\$8) and a box of candy (\$2). This means he will spend \$8 + \$2 = \$10.
Therefore, Colby will have \$20 - \$10 = \$10 left.
Now, let’s calculate how many movies he can see with the remaining \$10:
Since tickets on Fridays and Saturdays cost \$10, and he has \$10 left, he can see 1 movie on a Friday or Saturday.
Therefore, Colby can see 1 more movie on the last day of the month.
So, the final answer is 111.
@@
First, let’s calculate how many letters Mike receives in a week from the 5 pen pals. Each pen pal sends 2 letters per week, so 5 pen pals would send him 5 * 2 = 10 letters in a week.
Next, let’s calculate how many pages Mike needs to write in response to each letter. Each letter is 5 pages long, and Mike responds in kind, so he would write 5 pages for each letter he receives.
Therefore, Mike needs to write a total of 10 * 5 = 50 pages in a week.
Now, let’s calculate how many minutes it takes for Mike to write one page. We are given that he can write a page in 6 minutes.
Since Mike needs to write 50 pages, it will take him 50 * 6 = 300 minutes to complete all the writing.
Lastly, let’s convert the minutes to hours. There are 60 minutes in an hour, so 300 minutes is equal to 300 / 60 = 5 hours.
Therefore, Mike spends 5 hours writing in a week.
Answer: 555.
@@
First, let’s calculate the total number of letters that Mike receives in a week. Since he was pen pals with 5 people, and he stopped being pen pals with 2 of them, he is still receiving letters from 5 - 2 = 3 people.
Each of these 3 people sends 2 letters a week, so the total number of letters that Mike receives in a week is 3 * 2 = 6 letters.
Next, let’s calculate the total number of pages that Mike needs to write in response. Each letter is 5 pages long, so the total number of pages that Mike needs to write is 6 * 5 = 30 pages.
Since Mike can write a page every 6 minutes, the total time he spends writing in a week is 30 * 6 = 180 minutes.
To convert this to hours, we divide by 60 since there are 60 minutes in an hour. So, Mike spends 180 / 60 = 3 hours writing in a week.
Therefore, Mike spends 333 hours writing a week.
@@
First, we know that the body of the sweater takes 900 stitches to complete.
Next, we know that the collar takes a tenth of the number of stitches as the body. So, the collar takes 900 stitches / 10 = 90 stitches.
Finally, we know that the rosette takes twice as many stitches as the collar. So, the rosette takes 90 stitches * 2 = 180 stitches.
To find out how many stitches each sleeve takes, we need to subtract the stitches of the body, collar, and rosette from the total number of stitches in the sweater.
So, each sleeve takes 1800 stitches - 900 stitches (body) - 90 stitches (collar) - 180 stitches (rosette) = 630 stitches.
Therefore, each sleeve takes 630630630 stitches.
@@
First, we know that the body of the sweater takes 900 stitches to complete.
Next, we are told that the collar takes a tenth of the number of stitches in the body of the sweater. Since the body has 900 stitches, the collar will have 900 / 10 = 90 stitches.
Finally, we are told that the rosette takes twice as many stitches as the collar. Since the collar has 90 stitches, the rosette will have 90 * 2 = 180 stitches.
Now, let’s add up all the stitches to find the total number of stitches in the sweater: 900 (body) + 90 (collar) + 180 (rosette) = 1170 stitches.
We are given that the whole sweater is an 1800-stitch project. So, the remaining stitches must be for the sleeves.
1800 (total stitches) - 1170 (stitches accounted for) = 630 stitches for the sleeves.
Since there are two sleeves, each sleeve will have 630 / 2 = 315 stitches.
Therefore, each sleeve takes 315 stitches.
@@
First, let’s calculate how many pages Solo has left to read. He needs to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook, and 8 pages from his Geography textbook.
So, the total number of pages he needs to read is 4 + 20 + 7 + 8 = 39 pages.
Since Solo has 4 more days to complete all his reading, we can divide the total number of pages by the number of days.
Therefore, Solo needs to read an average of 39 pages / 4 days = 9.75 pages per day.
Therefore, Solo needs to read an average of 9.759.759.75 pages per day.
@@
First, let’s calculate the total number of pages Solo needs to read. He needs to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook, and 8 pages from his Geography textbook. So, the total number of pages he needs to read is 4 + 20 + 7 + 8 = 39 pages.
Next, let’s calculate the number of pages he has left to read after Monday. He read 15 pages on Monday, so he has 39 - 15 = 24 pages left to read.
Finally, let’s calculate the average number of pages he needs to read per day. He has 4 more days to complete his reading, so he needs to read 24 / 4 = 6 pages per day.
Therefore, Solo needs to read an average of 666 pages per day.
@@
First, let’s consider the size of a nickel. A nickel has a diameter of approximately 21.21 mm and a thickness of approximately 1.95 mm.
Next, let’s consider the size of a koala pouch. A koala pouch is designed to hold a baby koala, also known as a joey. The size of a koala pouch can vary, but it is generally not large enough to fit objects the size of a nickel.
Therefore, based on the size of a nickel and the size of a koala pouch, it is unlikely that a nickel would fit inside a koala pouch.
$𝑁𝑜$
@@
First, let’s consider the size of a nickel. A nickel has a diameter of about 0.835 inches (21.21 mm) and a thickness of about 0.077 inches (1.95 mm).
Next, let’s consider the size of a koala pouch. A koala pouch is designed to hold a baby koala, also known as a joey. The pouch is stretchy and expands as the joey grows. However, the size of the pouch may vary depending on the age and size of the joey.
Based on the size of a nickel and the stretchiness of a koala pouch, it is possible that a nickel could fit inside a koala pouch. However, it would depend on the size of the pouch and the position of the joey inside.
So, the answer is $𝑦𝑒𝑠$, a nickel could potentially fit inside a koala pouch.

### v3

@@ Abstract
Xiaohan Xu1thanks:   This work was done during internship at Microsoft., Chongyang Tao2, Tao Shen3, Can Xu2, Hongbo Xu1, Guodong Long3, Jian-guang Lou2 1Institute of Information Engineering, CAS, {xuxiaohan,hbxu}@iie.ac.cn 2Microsoft Corporation, {chotao,caxu,jlou}@microsoft.com 3AAII, School of CS, FEIT, UTS, {tao.shen,guodong.long}@uts.edu.au
@@ 2.2 Re-Reading (Re2) Improves Reasoning
Re2 Prompting Q: {Input Query} Read the question again: {Input Query} # Thought-eliciting prompt (e.g.,‘‘Let’s think step by step") #
@@ 3 Experiments
Table 1: Results on arithmetic reasoning benchmarks. denotes that Vanilla is even superior to CoT prompting.
@@
Table 2: Results on commonsense and symbolic reasoning benchmarks. denotes that Vanilla is even superior to CoT prompting.
@@ 3.3 Evaluation Results
When applied to ChatGPT, Re2 exhibits consistent improvement on most datasets, except for a slight drop in performance on a few datasets, e.g., AQUA and MultiArith, when using Vanilla+Re2. This exception could be due to ChatGPT’s exposure to these datasets with CoT outputs during instruction fine-tuning (IFT) Chen et al. (2023). On such datasets, ChatGPT with Vanilla still produces CoT-like output (see examples in Appendix E) and even outperforms ChatGPT with CoT (as indicated by the results in Tables 1 and 2). Chen et al. (2023) obtained similar experimental results and suggested that this occurs because ChatGPT may have been exposed to these task datasets containing CoT explanations without explicit prompting. Therefore, additional explicit instructions, like CoT or Re2, might disrupt this learned pattern in ChatGPT, possibly leading to decreased performance. Nonetheless, on some datasets like SVAMP, ASDIV, CSQA, and Date, Re2 still manages to improve the baseline Vanilla prompting. Moreover, in datasets where CoT prompting normally surpasses Vanilla prompting, such as GSM, StrategyQA, and Coin, Re2 significantly enhances Vanilla prompting ($\uparrow$ 4.63 on StrategyQA and $\uparrow$ 5.20 on the Coin dataset). Overall, our Re2 method still achieves improvements in 71% of the experiments on ChatGPT. More examples from the experiment results can be found in Appendix E.
@@ 3.4 Discussions
P0
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
77.79 78.77
P1
Q: {question}
Read the question again: {question}
#Answer format instruction#
A: Let’s think step by step.
79.45 80.59
P2
Q: {question}
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
78.09 79.38
P3
Q: {question}
A: Let’s think step by step.
Read the question again: {question}
#Answer format instruction#
A: Let’s think step by step.
79.08 80.36
P4
Q: {question}
A: Let’s think step by step.
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
78.09 79.38
@@ Appendix E More Cases
Vanilla
Q: {question}
#Answer format instruction#
A:
Vanilla+Re2
Q: {question}
Read the question again: {question}
#Answer format instruction#
A:
CoT
Q: {question}
#Answer format instruction#
A: Let’s think step by step.
CoT+Re2
Q: {question}
Read the question again: {question}
#Answer format instruction#
A: Let’s think step by step.
PS
Q: {question}
#Answer format instruction#
A: Let’s first understand the problem and devise a plan to solve the problem.
Then, let’s carry out the plan, solve the problem step by step,
and give the ultimate answer. Please explicitly generate the mentioned process:
[Problem Understanding], [Plan], [Solving/Calculations], [Answer]. in your response.
PS+Re2
Q: {question}
Read the question again: {question}
#Answer format instruction#
A: Let’s first understand the problem and devise a plan to solve the problem.
Then, let’s carry out the plan, solve the problem step by step,
and give the ultimate answer. Please explicitly generate the mentioned process:
[Problem Understanding], [Plan], [Solving/Calculations], [Answer]. in your response.
@@
PAL
#!/bin/python3
import math
import numpy as np
import statistics
import sympy as sp
########### Task Instruction ##############
# You will write python program to solve math problems.
# You will only write code blocks.
# Please generate your code block in ‘def solution()‘ function, and thus it can be
executed by python interpreter. You don’t need to call ‘solution()‘ function because
it will be called by the system.
# The concrete format of ‘solution()‘ is as follows:
# def solution():
# """<question>"""
# <your code>
# result = <your result>
# return result
# Q: {question}
# Your defined "solution()" function with comments here.
PAL+Re2
#!/bin/python3
import math
import numpy as np
import statistics
import sympy as sp
########### Task Instruction ##############
# You will write python program to solve math problems.
# You will only write code blocks.
# Please generate your code block in ‘def solution()‘ function, and thus it can be
executed by python interpreter. You don’t need to call ‘solution()‘ function because
it will be called by the system.
# The concrete format of ‘solution()‘ is as follows:
# def solution():
# """<question>"""
# <your code>
# result = <your result>
# return result
# Q: {question}
# Read the question again: {question}
# Your defined "solution()" function with comments here.
@@
GSM
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
SVAMP
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
ASDIV
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
AQUA
@@
MultiArith
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
SingleEQ
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
AddSub
Your final answer should be a single numerical number, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$,
at the end of your response.
CommonsenseQA
@@
StrategyQA
Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
ARC
@@
Date Understanding
Your answer should be a date, in the format of $𝑀𝑀𝐷𝐷𝑌𝑌𝑌𝑌$,
e.g. 05/01/20220501202205 / 01 / 2022.
Coin Flip
Your answer should be either $𝑦𝑒𝑠$ or $𝑛𝑜$, in the form $𝑎𝑛𝑠𝑤𝑒𝑟$.
@@
First, let’s calculate how much feed Wendi gives her chickens in the morning. She has 20 chickens, and she gives each chicken 3 cups of feed, so in total, she gives them 20 x 3 = 60 cups of feed in the morning. Next, let’s calculate how much feed Wendi gives her chickens in the afternoon. Again, she has 20 chickens, and she gives each chicken 3 cups of feed, so in total, she gives them 20 x 3 = 60 cups of feed in the afternoon. To find out how much feed she needs to give her chickens in the final meal of the day, we need to subtract the total feed given in the morning and afternoon from the total feed needed for the day. The total feed needed for the day for 20 chickens is 20 x 3 x 3 = 180 cups of feed. So, the feed needed in the final meal of the day is 180 - 60 - 60 = 60 cups of feed. Therefore, Wendi needs to give her chickens 60 cups of feed in the final meal of the day. Answer: 606060
@@
First, we need to determine the total amount of feed Wendi gives her chickens in the morning and afternoon. In the morning, she gives 15 cups of feed, and in the afternoon, she gives another 25 cups. So, the total amount of feed given in the morning and afternoon is 15 + 25 = 40 cups. Next, we need to determine the total amount of feed needed for Wendi’s flock of 20 chickens. Since each chicken requires 3 cups of feed, we multiply 3 by 20 to get 60 cups. Finally, we subtract the total amount of feed given in the morning and afternoon from the total amount needed. 60 - 40 = 20 Therefore, Wendi needs to give her chickens 20 cups of feed in the final meal of the day. Answer: 202020
@@
First, let’s calculate the cost of the movies Colby has already seen. He has seen 5 movies on a Friday or Saturday, which cost \$10 each, so that is a total of 5 * \$10 = \$50. He has also seen 8 movies on other days, which cost \$7 each, so that is a total of 8 * \$7 = \$56. Therefore, the total cost of the movies he has already seen is \$50 + \$56 = \$106. Next, let’s calculate the cost of the popcorn and candy he wants to buy on the last day of the month. The popcorn costs \$8 and the candy costs \$2, so the total cost of the snacks is \$8 + \$2 = \$10. Now, let’s subtract the cost of the movies Colby has already seen and the cost of the snacks from the \$150 his parents gave him. \$150 - \$106 - \$10 = \$34. Since each movie on a Friday or Saturday costs \$10 and Colby has \$34 left, he can see a maximum of \$34 / \$10 = 3 movies on a Friday or Saturday. In conclusion, Colby can see a total of 3 + 5 = 8 movies in the month. 888
@@
First, let’s calculate the total amount Colby has spent so far: For the 5 movies on Fridays and Saturdays, he spent 5 x \$10 = \$50. For the 8 movies on other days, he spent 8 x \$7 = \$56. For the 2 tubs of popcorn, he spent 2 x \$8 = \$16. For the 4 boxes of candy, he spent 4 x \$2 = \$8. Adding all these expenses together, Colby has spent \$50 + \$56 + \$16 + \$8 = \$130. Now, let’s calculate how much money he has left from the \$150 his parents gave him: Colby started with \$150 - \$130 = \$20. Since it’s the last day of the month and it’s a Friday, he wants to buy a popcorn (\$8) and a box of candy (\$2). This means he will spend \$8 + \$2 = \$10. Therefore, Colby will have \$20 - \$10 = \$10 left. Now, let’s calculate how many movies he can see with the remaining \$10: Since tickets on Fridays and Saturdays cost \$10, and he has \$10 left, he can see 1 movie on a Friday or Saturday. Therefore, Colby can see 1 more movie on the last day of the month. So, the final answer is 111.
@@
First, let’s calculate how many letters Mike receives in a week from the 5 pen pals. Each pen pal sends 2 letters per week, so 5 pen pals would send him 5 * 2 = 10 letters in a week. Next, let’s calculate how many pages Mike needs to write in response to each letter. Each letter is 5 pages long, and Mike responds in kind, so he would write 5 pages for each letter he receives. Therefore, Mike needs to write a total of 10 * 5 = 50 pages in a week. Now, let’s calculate how many minutes it takes for Mike to write one page. We are given that he can write a page in 6 minutes. Since Mike needs to write 50 pages, it will take him 50 * 6 = 300 minutes to complete all the writing. Lastly, let’s convert the minutes to hours. There are 60 minutes in an hour, so 300 minutes is equal to 300 / 60 = 5 hours. Therefore, Mike spends 5 hours writing in a week. Answer: 555.
@@
First, let’s calculate the total number of letters that Mike receives in a week. Since he was pen pals with 5 people, and he stopped being pen pals with 2 of them, he is still receiving letters from 5 - 2 = 3 people. Each of these 3 people sends 2 letters a week, so the total number of letters that Mike receives in a week is 3 * 2 = 6 letters. Next, let’s calculate the total number of pages that Mike needs to write in response. Each letter is 5 pages long, so the total number of pages that Mike needs to write is 6 * 5 = 30 pages. Since Mike can write a page every 6 minutes, the total time he spends writing in a week is 30 * 6 = 180 minutes. To convert this to hours, we divide by 60 since there are 60 minutes in an hour. So, Mike spends 180 / 60 = 3 hours writing in a week. Therefore, Mike spends 333 hours writing a week.
@@
First, we know that the body of the sweater takes 900 stitches to complete. Next, we know that the collar takes a tenth of the number of stitches as the body. So, the collar takes 900 stitches / 10 = 90 stitches. Finally, we know that the rosette takes twice as many stitches as the collar. So, the rosette takes 90 stitches * 2 = 180 stitches. To find out how many stitches each sleeve takes, we need to subtract the stitches of the body, collar, and rosette from the total number of stitches in the sweater. So, each sleeve takes 1800 stitches - 900 stitches (body) - 90 stitches (collar) - 180 stitches (rosette) = 630 stitches. Therefore, each sleeve takes 630630630 stitches.
@@
First, we know that the body of the sweater takes 900 stitches to complete. Next, we are told that the collar takes a tenth of the number of stitches in the body of the sweater. Since the body has 900 stitches, the collar will have 900 / 10 = 90 stitches. Finally, we are told that the rosette takes twice as many stitches as the collar. Since the collar has 90 stitches, the rosette will have 90 * 2 = 180 stitches. Now, let’s add up all the stitches to find the total number of stitches in the sweater: 900 (body) + 90 (collar) + 180 (rosette) = 1170 stitches. We are given that the whole sweater is an 1800-stitch project. So, the remaining stitches must be for the sleeves. 1800 (total stitches) - 1170 (stitches accounted for) = 630 stitches for the sleeves. Since there are two sleeves, each sleeve will have 630 / 2 = 315 stitches. Therefore, each sleeve takes 315 stitches.
@@
First, let’s calculate how many pages Solo has left to read. He needs to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook, and 8 pages from his Geography textbook. So, the total number of pages he needs to read is 4 + 20 + 7 + 8 = 39 pages. Since Solo has 4 more days to complete all his reading, we can divide the total number of pages by the number of days. Therefore, Solo needs to read an average of 39 pages / 4 days = 9.75 pages per day. Therefore, Solo needs to read an average of 9.759.759.75 pages per day.
@@
First, let’s calculate the total number of pages Solo needs to read. He needs to read 4 pages from his Science textbook, 20 pages from his Social Studies textbook, 7 pages from his History textbook, and 8 pages from his Geography textbook. So, the total number of pages he needs to read is 4 + 20 + 7 + 8 = 39 pages. Next, let’s calculate the number of pages he has left to read after Monday. He read 15 pages on Monday, so he has 39 - 15 = 24 pages left to read. Finally, let’s calculate the average number of pages he needs to read per day. He has 4 more days to complete his reading, so he needs to read 24 / 4 = 6 pages per day. Therefore, Solo needs to read an average of 666 pages per day.
@@
First, let’s consider the size of a nickel. A nickel has a diameter of approximately 21.21 mm and a thickness of approximately 1.95 mm. Next, let’s consider the size of a koala pouch. A koala pouch is designed to hold a baby koala, also known as a joey. The size of a koala pouch can vary, but it is generally not large enough to fit objects the size of a nickel. Therefore, based on the size of a nickel and the size of a koala pouch, it is unlikely that a nickel would fit inside a koala pouch. $𝑁𝑜$
@@
First, let’s consider the size of a nickel. A nickel has a diameter of about 0.835 inches (21.21 mm) and a thickness of about 0.077 inches (1.95 mm). Next, let’s consider the size of a koala pouch. A koala pouch is designed to hold a baby koala, also known as a joey. The pouch is stretchy and expands as the joey grows. However, the size of the pouch may vary depending on the age and size of the joey. Based on the size of a nickel and the stretchiness of a koala pouch, it is possible that a nickel could fit inside a koala pouch. However, it would depend on the size of the pouch and the position of the joey inside. So, the answer is $𝑦𝑒𝑠$, a nickel could potentially fit inside a koala pouch.